OpenStudy (anonymous):

HELP ! the area of the kite is 36 cm squared .. the length of the diagonal SO is 2x and the diagonal PI is 3x... what is the value of x?

OpenStudy (anonymous):



OpenStudy (anonymous):

area of kite = area of triangle PSL + area of triangle POL

OpenStudy (campbell_st):

just use Area = 1/2 * product of the diagonals... so 36 = 1/2*(3x)*(2x)

OpenStudy (anonymous):

Area of triangle PML \[=\frac{1}{2}SM*PL\] \[=\frac{1}{2}x*3x\] \[=\frac{3x^{2}}{2}\] Area of triangle POL \[=\frac{1}{2}OM*PL\] \[=\frac{1}{2}x*3x\] \[=\frac{3x^{2}}{2}\] so , \[36=\frac{3x^2}{2}+\frac{3x^2}{2}\] Now can you solve?

OpenStudy (anonymous):

are you saying that x=3 ??

OpenStudy (anonymous):

@campbell_st I tried but it didn't work

OpenStudy (campbell_st):

so \[36 = \frac{6x^2}{2} \] which becomes \[36 = 3x^2\] divide both sides by 3 then that the square root of both sides to find x